# Hierarchical Concept Embedding & Pursuit for Interpretable Image Classification

### v1

@@ 3 Hierarchical Concept Embedding Model
Figure 3: An example hierarchy with $L=2$ levels, branching factor $b=2$, and $N_{L}=6$ nodes in total. Each node $i\in\mathcal{A}$ has an associated vector representation $\bm{a}^{(i)}\in\mathbb{R}^{d}$, where $\mathcal{A}=\{1,\ldots,N_{L}\}$ is the set of node indices. Let $\operatorname{par}(\cdot),\operatorname{chi}(\cdot),\operatorname{anc}(\cdot),\operatorname{desc}(\cdot):\mathcal{A}\rightarrow\mathcal{P}(\mathcal{A})$ respectively be functions returning the set of parent, children, ancestors, and descendants of a node. The level of a node is the number of its ancestors, that is $\operatorname{lev}(i)=|\operatorname{anc}(i)|$.
@@ 3.1 Well-clustered synset embeddings
To ensure that each node in the hierarchy can be uniquely assigned to its parent, we impose geometric constraints on the embedding space. The following proposition formalizes conditions that guarantee well-separated subtrees and unambiguous parent identification.
@@
All proofs are provided in App. B. A sufficient way to satisfy both conditions in Proposition 3.1 is through a geometric half-angle decreasing schedule, as presented next.
@@ 4.1 Hierarchical Dictionary Construction
$\small\!\!\!\!\Bigl[\,\bm{a}^{(1)},\ldots,\bm{a}^{(b)},\;\bm{a}^{(j_{1})}-\bm{a}^{(\operatorname{par}(j_{1}))},\ldots,\bm{a}^{(j_{k})}-\bm{a}^{(\operatorname{par}(j_{k}))}\,\Bigr]\!\!$ (5)
where $\left(j_{1},\ldots,j_{k}\right)=\left(j\in\mathcal{A}:j>b\right)$. The first $b$ columns contain the root child atoms, while the remaining columns contain the differences between synsets and their parents for non-root nodes.
In the context of interpretable image classification, $\bm{a}^{(i)}$ represents the embedding of synsets within the WordNet [50] hierarchy. [^4] We use WordNet as a concrete example, but this analysis generalizes to other hierarchies. while the difference $\bm{a}^{(i)}-\bm{a}^{(\operatorname{par}(i))}$ is the embedding of concepts that distinguish the synset from its parent. As an example, if $\bm{a}^{(i)}$ is the embedding of the synset bear and $\bm{a}^{(j)}-\bm{a}^{(i)}$ is the embedding of the concept the color white, then $\bm{a}^{(j)}$ is the embedding of the synset polar bear. This construction of the dictionary in Eq. 5 avoids the trivial solution to the sparse formulation of concept recovery (i.e., the sparsest explanation for an image of a cat is just that it is a cat) Note that the difference embeddings have a grounded and interpretable meaning on their own, as they represent the direction that differentiates a child synset from its parent.
With this model, we can verify that a synset embedding is a sparse linear combination of the atoms corresponding to the nodes along the path from the root to the leaf node $i$
$\bm{a}^{(i)}=\bm{a}^{(\pi_{1}(i))}+\sum_{\begin{subarray}{c}j\in\operatorname{anc}(i);~j\neq\pi_{1}(i)\end{subarray}}\bigl(\bm{a}^{(j)}-\bm{a}^{(\operatorname{par}(j))}\bigr),$ (6)
for some synset $i\in\mathcal{A}$ and $\pi_{1}(i)$ denotes the first ancestor of $i$. If we assume that an image embedding $\bm{x}$ is generated from the synset $i$ with some small noise, i.e., $\bm{x}=\bm{a}^{(i)}+\epsilon$, recovering the synset $i$ is equivalent to recovering the sparse code $\bm{z}$ with non-zero entries corresponding to the nodes along the path from the root to $i$. An illustration of this hierarchical sparse decomposition is shown in Fig. 4.
@@ 4.2 Hierarchical Orthogonal Matching Pursuit
2:Init with a null hypothesis set: $\mathcal{H}^{(0)}\leftarrow\Big\{\big(\emptyset,\bm{x},\bm{0}\big)\Big\}$
@@
8:   $\mathcal{D}_{\text{active}}\leftarrow\textsc{ExtendDict}(h,t)$ (Alg. 2)
@@
1:iteration $t$, hypothesis $h=(\mathcal{S},\bm{r},i_{last})$, roots $\mathcal{R}$
2:if $t=0$ then $\triangleright$ root children at first step
3:  $\textbf{return}\;\bm{D}_{\text{active}}=\{\bm{a}_{1,i}:i\in\operatorname{chi}(\mathcal{R})\}$
@@
There are two main modifications from OMP: (1) Extending the sparse support only by children of the deepest node explored; (2) Using beam search to manage the hypothesis set. The overall procedure is detailed in Algorithm 1. (For completeness, we describe OMP in Algorithm 5). Note that for clarity, which level a synset is at, a synset embedding $\bm{a}^{(i)}$ in the algorithm is referred to as $\bm{a}_{l,i}$ where $l$ is the depth of node $i$. Also, except for the root children atoms, all other atoms are not the synset themselves but the differences of the synsets from their parents. We formalize the theoretical advantages of this hierarchical approach in the following proposition.
@@ 5.1 Synthetic Experiments
We choose a branching factor of $b=3$, hierarchy depth $L=7$, and dimension $d=50$. Note that this dimension satisfies the depth–dimension condition in Eq. 37 since $d=50\geq 7+3=10$. This gives us $2187$ leaf synsets and $3280$ atoms in the dictionary. We generate 5 samples per leaf for a total of $10,935$ samples. Further details on how the synthetic data is generated can be found in App. C, and the hyperparameters can be found in Sec. D.1.
@@ 5.2 Real-world Experiments
In this section, we evaluate HCEP on real-world image classification tasks. First, given a class (which is a leaf node), we estimate the class embeddings by taking the mean of the CLIP [56] image embeddings for images belonging to that class. For the non-leaf synsets, we estimate their embeddings using the mean of their children’s embeddings. Next, we construct the hierarchical dictionary as described in Eq. (5). We keep this fixed dictionary for all experiments. We next provide the overall experiment settings; more details are in Sec. D.2.
@@
Figure 6: Observed Hierarchical Orthogonality on ImageNet for CLIP [56]. The cosine similarity between child-parent difference vectors and their parents is close to zero, while random non-parent pairs have significantly higher cosine similarity. This suggests that hierarchical orthogonality [52] holds even on contrastively trained vision models.
@@
Table 1: As we reduce the number of images per class in ImageNet, HCEP consistently improves test classification accuracy, support precision, and support recall over all baselines. Result at sparsity level 14.
Method Classification Accuracy Support Precision Support Recall
12-shot 25-shot 50-shot Full 12-shot 25-shot 50-shot Full 12-shot 25-shot 50-shot Full
OMP 45.8 52.9 57.4 70.7 15.6 16.3 16.6 17.5 36.6 37.9 38.3 41.0
Hierarch. NN 28.1 29.0 29.8 15.7 47.9 48.2 48.6 34.9 44.1 44.0 44.1 28.7
CBM 24.7 34.4 44.0 78.5 30.5 31.8 33.3 45.6 60.0 62.9 66.2 97.3
HCEP (Ours) 52.1 57.3 61.3 65.2 71.2 72.2 73.0 70.1 71.8 72.7 73.4 71.1
@@
See the results on ImageNet (Tab. 1), CIFAR100/ImageNette (Fig. 8). We observe that Hierarchical OMP achieves higher support precision and recall compared to other baselines, indicating better recovery of relevant concepts. In the low-data setting, Hierarchical OMP outperforms all interpretable baselines in classification accuracy and support precision/recall, demonstrating its robustness in low-data settings.
We also verified that HCEP generalizes across vision-language models: replacing CLIP with SigLIP [74] on ImageNet yields similar improvements in interpretability metrics (see Fig. 11).
Runtime Analysis. Fig. 9 shows a runtime analysis using the same setting as in Tab. 1. As HCEP’s beam width increases from $1$ to $32$, HCEP achieves stronger support precision with a modest runtime overhead, achieved by parallelizing the beam search hypotheses on GPU.
@@ A.3 Sparse Recovery
We briefly recall two classical sparse recovery approaches that motivate our hierarchical construction: greedy pursuit via Orthogonal Matching Pursuit (OMP) and convex relaxation via Basis Pursuit (BP).
@@
with normalized columns (atoms) of $\bm{D}$. We focus on the exact linear model for clarity. The goal is to recover the unknown sparse coefficient vector $\bm{z}$ from $\bm{x}$. We use the shorthand $k]:=\{1,2,\ldots,k\}$ for column indices.
@@
The classical MP algorithm greedily selects at each step the dictionary atom most correlated with the current residual and updates the residual by removing its component along that atom, without re-fitting over the accumulated support [47]. OMP below is the orthogonalized variant that re-solves on the active set at every iteration.
Orthogonal Matching Pursuit (OMP).
OMP is a greedy algorithm that orthogonalizes over the active set at each step (See OMP in Algorithm 5; for comparison, we also provide the full details of Hierarchical OMP in Algorithm 6), contrasting with our hierarchical variant. The restricted least-squares update used at each iteration is
@@
Appendix B Proofs
B.1 Proof of Proposition 3.1
@@ Appendix B Proofs
Subtree containment (Eq. (2)): Since $k\in\operatorname{desc}(j)$, Eq. (2) gives
$\angle(\bm{a}^{(j)},\bm{a}^{(k)})\leq\theta_{\operatorname{lev}(j)}.$ (12)
Thus, every descendant of $j$ lies within the cone of half-angle $\theta_{\operatorname{lev}(j)}$ rooted at $\bm{a}^{(j)}$, so in particular $k$ is confined to this cone.
Sibling-cone disjointness (Eq. (3)): Consider any sibling $j^{\prime}\in\operatorname{chi}(i)$ with $j^{\prime}\neq j$. To derive a contradiction, assume that $k$ also lies in the subtree of $j^{\prime}$, i.e., $k\in\operatorname{desc}(j^{\prime})$. Then, by subtree containment applied to $j^{\prime}$, we similarly obtain
@@
Consider the spherical triangle formed by the unit vectors $\bm{a}^{(j)}/\|\bm{a}^{(j)}\|$, $\bm{a}^{(k)}/\|\bm{a}^{(k)}\|$, and $\bm{a}^{(j^{\prime})}/\|\bm{a}^{(j^{\prime})}\|$ on the unit sphere. By the spherical triangle inequality, the angle between any two vertices is at most the sum of the angles to the third vertex:
@@
B.2 Proof of Proposition 3.2
Statement. If the half-angles satisfy the geometric decrease $\theta_{l+1}\leq r\,\theta_{l}$ with $r\in(0,1/2)$, then subtree containment (Eq. (2)) holds. Moreover, a simple sufficient condition to realize sibling-cone disjointness (Eq. (3)) together with packing of $b$ sibling cones under a parent cone of half-angle $\theta_{l}$ is
$\theta_{l+1}\leq\frac{1}{b}\,\theta_{l}.$ (15)
Consequently, if the half-angles satisfy $\theta_{l+1}\leq\min\{r,1/b\}\,\theta_{l}$ with $r\in(0,1/2)$, then there exists a placement of the nodes such that subtree containment (Eq. (2)) and sibling-cone disjointness (Eq. (3)) hold.
@@
which proves Eq. (16) and hence the subtree cone condition in Eq. (2).
To see that Eq. (15) is sufficient, consider any 2D plane containing the parent cone axis. In this plane, a cone of half-angle $\alpha$ appears as a planar angle of magnitude $2\alpha$. Placing the $b$ child cone axes within this plane and packing $b$ child angles $2\theta_{l+1}$ inside the parent angle $2\theta_{l}$ is possible whenever $b\theta_{l+1}\leq\theta_{l}$, i.e. Eq. (15).
B.3 Proof of Proposition 3.3
@@
B.4 Intermediate results for Proposition 4.1
Lemma B.1 (Column normalization equivalence).
@@
Immediate from $\widehat{\bm{d}}_{j}=\bm{d}_{j}/\|\bm{d}_{j}\|_{2}\,$ and $\langle\bm{r},\widehat{\bm{d}}_{j}\rangle=\langle\bm{r},\bm{d}_{j}\rangle/\|\bm{d}_{j}\|_{2}$, together with $\bm{x}=\bm{D}\bm{z}=\bm{D}\,\bm{W}^{-1}\bm{W}\bm{z}=\widehat{\bm{D}}\,\widehat{\bm{z}}$. Since diagonal rescaling of the columns in $\bm{D}_{S}$ leaves the column span unchanged, the orthogonal projector onto $\mathrm{span}(\bm{D}_{S})$ is invariant to such rescaling. Therefore, once the same index is selected, the least-squares updates utilize the same projector, and the residuals match at every step. ∎
Definition B.2 (ERC on normalized dictionary).
 [^7] This follows the classical exact recovery coefficient (ERC) of Tropp [66].
@@
$\mathrm{ERC}(\widehat{\bm{D}};S)$ $:=\big\|\,\widehat{\bm{D}}_{S}^{\dagger}\,\widehat{\bm{D}}_{S^{c}}\,\big\|_{\infty}$, (23)
$\mathrm{ERC}(\widehat{\bm{D}};S\,|\,T)$ $:=\big\|\,\widehat{\bm{D}}_{S}^{\dagger}\,\widehat{\bm{D}}_{T\setminus S}\,\big\|_{\infty}$, (24)
for any $T\supseteq S$.
Lemma B.3 (Monotone ERC improvement under subtree restriction).
@@
The quantity $\widehat{\bm{D}}_{S}^{\dagger}$ is fixed, and shrinking $T$ only removes columns from $\widehat{\bm{D}}_{T\setminus S}$, so the maximum defining the ERC is taken over a subset and therefore cannot increase. ∎
Lemma B.4 (ERC threshold implies restricted OMP success).
Under the assumptions of Lemma B.3, if $\mathrm{ERC}(\widehat{\bm{D}};S\,|\,T_{L})<1$, then OMP run on $\bm{D}$ with the normalized selection rule
@@
Lemma B.1 shows that the normalized-selection rule on $\bm{D}$ matches standard OMP on $\widehat{\bm{D}}$. The classical noiseless ERC theorem of Tropp [66] applied to the restricted dictionary $\widehat{\bm{D}}_{T_{L}}$ then yields exact support recovery in $s$ iterations whenever $\|\widehat{\bm{D}}_{S}^{\dagger}\,\widehat{\bm{D}}_{T_{L}\setminus S}\|_{\infty}<1$. ∎
B.5 Proof of Proposition 4.1
@@
If the maximizer(s) contributing to $\mathrm{ERC}(\widehat{\bm{D}};S)$ lie outside $T_{L}$, pruning them ensures $\mathrm{ERC}(\widehat{\bm{D}};S\,|\,T_{L})<\mathrm{ERC}(\widehat{\bm{D}};S)$, so the restricted value can fall below $1$ while the global one remains at least $1$. Whenever this happens, Lemma B.4 certifies exact recovery for Hierarchical OMP on $T_{L}$, whereas the ERC test for OMP on the full dictionary fails. Thus, the subtree-restricted algorithm possesses a strictly larger guaranteed support-recovery region. ∎
Appendix C Step-by-step construction of a Hierarchical Concept Embedding

### v2

@@ 3 Hierarchical Concept Embedding Model
Figure 3: A hierarchy with $L=2$ levels, branching factor $b=2$, and $N_{L}=6$ nodes in total. Each node $i\in\mathcal{A}$ has an associated vector representation $\bm{a}^{(i)}\in\mathbb{R}^{d}$, where $\mathcal{A}=\{1,\ldots,N_{L}\}$ is the set of non-root node indices. Let $\operatorname{par}(\cdot),\operatorname{chi}(\cdot),\operatorname{anc}(\cdot),\operatorname{desc}(\cdot)$ respectively be functions returning the set of parent, children, ancestors, and descendants of a node. The level of a node is the number of its ancestors, that is $\operatorname{lev}(i)=|\operatorname{anc}(i)|$.
@@ 3.1 Well-clustered synset embeddings
To ensure that each node in the hierarchy can be uniquely assigned to its parent, we impose geometric constraints on the embedding space. The following proposition formalizes sufficient conditions that guarantee well-separated subtrees.
@@
All proofs are provided in App. C. A sufficient way to satisfy both conditions in Proposition 3.1 is through a geometric half-angle decreasing schedule, as presented next.
@@ 4.1 Hierarchical Dictionary Construction
$\bm{D}=\bigl[\,\bm{a}^{(j)}-\bm{a}^{(\operatorname{par}(j))}\,\bigr]_{j\in\mathcal{A}}.$ (5)
Each column captures the difference between a synset and its parent. We define $\bm{a}^{(\mathrm{root})}=\mathbf{0}$, so the atoms for root children reduce to $\bm{a}^{(j)}-\mathbf{0}=\bm{a}^{(j)}$.
In the context of interpretable image classification, $\bm{a}^{(i)}$ represents the embedding of synsets within the WordNet [50] hierarchy. [^4] We use WordNet as a concrete example, but this analysis generalizes to other hierarchies. while the difference $\bm{a}^{(i)}-\bm{a}^{(\operatorname{par}(i))}$ is the embedding of concepts that distinguish the synset from its parent. As an example, if $\bm{a}^{(i)}$ is the embedding of the synset bear and $\bm{a}^{(j)}-\bm{a}^{(i)}$ is the embedding of the concept the color white, then $\bm{a}^{(j)}$ is the embedding of the synset polar bear. This construction of the dictionary in Eq. 5 avoids the trivial solution to the sparse formulation of concept recovery (i.e., the sparsest explanation for an image of a cat is just that it is a cat). Note that the difference embeddings have a grounded and interpretable meaning on their own, as they represent the direction that differentiates a child synset from its parent.
With this model, a synset embedding is a sparse linear combination of the atoms along the root-to-$i$ path:
$\bm{a}^{(i)}=\sum_{j\in\operatorname{anc}(i)\cup\{i\}}\bigl(\bm{a}^{(j)}-\bm{a}^{(\operatorname{par}(j))}\bigr),$ (6)
which follows by telescoping, since $\bm{a}^{(\mathrm{root})}=\mathbf{0}$. If we assume that an image embedding $\bm{x}$ is generated from the synset $i$ with some small noise, i.e., $\bm{x}=\bm{a}^{(i)}+\epsilon$, recovering the synset $i$ is equivalent to recovering the sparse code $\bm{z}$ with non-zero entries corresponding to the nodes along the path from the root to $i$. An illustration of this hierarchical sparse decomposition is shown in Fig. 4.
@@ 4.2 Hierarchical Orthogonal Matching Pursuit
2:Init with a null hypothesis set: $\mathcal{H}^{(0)}\leftarrow\Big\{\big(\emptyset,\bm{x},\mathrm{root}\big)\Big\}$
@@
8:   $\mathcal{D}_{\text{active}}\leftarrow\textsc{ExtendDict}(h)$ (Alg. 2)
@@
1:hypothesis $h=(\mathcal{S},\bm{r},i_{last})$
@@
There are two main modifications from OMP: (1) Extending the sparse support only by children of the deepest node explored; (2) Using beam search to manage the hypothesis set. The overall procedure is detailed in Algorithm 1. (For completeness, we describe OMP in Algorithm 5). We formalize the theoretical advantages of this hierarchical approach in the following proposition.
@@ 5.1 Synthetic Experiments
We choose a branching factor of $b=3$, hierarchy depth $L=7$, and dimension $d=50$. Note that this dimension satisfies the depth–dimension condition in Eq. 37 since $d=50\geq 7+3=10$. This gives us $2187$ leaf synsets and $3280$ atoms in the dictionary. We generate 5 samples per leaf for a total of $10,935$ samples. Further details on how the synthetic data is generated can be found in App. D, and the hyperparameters can be found in Sec. E.1.
@@
Figure 6: Observed Hierarchical Orthogonality on ImageNet for CLIP. The cosine similarity between child-parent difference vectors and their parents is close to zero, while random pairing has a significantly higher cosine similarity. This suggests that hierarchical orthogonality [53] holds even on pretrained models.
@@ 5.2 Real-world Experiments
In this section, we evaluate HCEP on real-world image classification tasks. First, given a class (which is a leaf node), we estimate the class embeddings by taking the mean of the CLIP [57] image embeddings for images belonging to that class. For the non-leaf synsets, we estimate their embeddings using the mean of their children’s embeddings. Next, we construct the hierarchical dictionary as described in Eq. (5). We keep this fixed dictionary for all experiments. We next provide the overall experiment settings; more details are in Sec. E.3.
@@
Figure 9: As we reduce the number of images per class in ImageNet, HCEP consistently improves test classification accuracy, support precision, and support recall over all baselines. Result at sparsity level 14.
@@
See the results on ImageNet (Fig. 9), CIFAR100/ImageNette (Fig. 8). We observe that Hierarchical OMP achieves higher support precision and recall compared to other baselines, indicating better recovery of relevant concepts. In the low-data setting, Hierarchical OMP outperforms all interpretable baselines in classification accuracy and support precision/recall, demonstrating its robustness in low-data settings. We also verified that HCEP generalizes across vision-language models: replacing CLIP with SigLIP [75] on ImageNet yields similar improvements in interpretability metrics (see Fig. 12).
Runtime Analysis. Fig. 10 shows a runtime analysis using the same setting as in Fig. 9. As HCEP’s beam width increases from $1$ to $32$, HCEP achieves stronger support precision with a modest runtime overhead, achieved by parallelizing the beam search hypotheses on GPU.
@@ A.3 Sparse Recovery
We briefly recall classical sparse recovery approaches that motivate our hierarchical construction.
@@
with normalized columns (atoms) of $\bm{D}$. We use the shorthand $k]:=\{1,2,\ldots,k\}$ for column indices. We focus on the exact linear model for clarity. The goal is to recover the unknown sparse coefficient vector $\bm{z}$ from $\bm{x}$, which is equivalent to the following $\ell_{0}$-minimization problem:
$\min_{\bm{z}\in\mathbb{R}^{k}}\|\bm{z}\|_{0}\quad\text{s.t.}\quad\bm{D}\bm{z}=\bm{x}.$ (9)
There are two strategies to solve this $\ell_{0}$ problem: greedy through matching pursuit and convex relaxation through basis pursuit.
@@
The classical MP algorithm greedily selects at each step the dictionary atom most correlated with the current residual and updates the residual by removing its component along that atom, without re-fitting over the accumulated support [47]. OMP is the orthogonalized variant that re-solves on the active set at every iteration (See OMP in Algorithm 5; for comparison, we also provide an expanded version of Hierarchical OMP in Algorithm 6), contrasting with our hierarchical variant. The restricted least-squares update used at each iteration is
@@
Appendix B Limitations
Appendix C Proofs
C.1 Proof of Proposition 3.1
@@ Appendix C Proofs
Eq. (2) states that the angle between $\bm{a}^{(j)}$ and $\bm{a}^{(k)}$ is at most $\theta_{\operatorname{lev}(j)}$, so the position of $k$ is constrained within the cone of half-angle $\theta_{\operatorname{lev}(j)}$ rooted at $\bm{a}^{(j)}$.
Consider any sibling $j^{\prime}\neq j$. To derive a contradiction, assume that $k$ also lies in the subtree of $j^{\prime}$, i.e., $k\in\operatorname{desc}(j^{\prime})$. Then, by subtree containment (Eq. (2)) applied to $j^{\prime}$, we similarly obtain
@@
Consider the spherical triangle formed by the unit vectors $\bm{a}^{(j)}/\|\bm{a}^{(j)}\|$, $\bm{a}^{(k)}/\|\bm{a}^{(k)}\|$, and $\bm{a}^{(j^{\prime})}/\|\bm{a}^{(j^{\prime})}\|$ on the unit sphere. By the triangle inequality, we have that:
@@
C.2 Proof of Proposition 3.2
Statement.
If the half-angles satisfy $\theta_{l+1}\leq\min\{r,1/b\}\,\theta_{l}$ with $r\in(0,1/2)$, then there exists a placement of the nodes such that subtree containment (Eq. (2)) and sibling-cone disjointness (Eq. (3)) hold.
@@
which proves Eq. (15) and hence the subtree containment condition in Eq. (2).
Next we show that the sibling-cone disjointness condition (Eq. (3)) can be satisfied by a simple sufficient condition on the half-angles:
$\theta_{l+1}\leq\frac{1}{b}\,\theta_{l}.$ (21)
Consider any 2D plane containing the parent cone axis. In this plane, a cone of half-angle $\alpha$ appears as a planar angle of magnitude $2\alpha$. Placing the $b$ child cone axes within this plane and packing $b$ child angles $2\theta_{l+1}$ inside the parent angle $2\theta_{l}$ is possible whenever $b\theta_{l+1}\leq\theta_{l}$, i.e. Eq. (21).
C.3 Proof of Proposition 3.3
@@
C.4 Intermediate results for Proposition 4.1
Lemma C.1 (Column normalization equivalence).
@@
Immediate from $\widehat{\bm{d}}_{j}=\bm{d}_{j}/\|\bm{d}_{j}\|_{2}\,$ and $\langle\bm{r},\widehat{\bm{d}}_{j}\rangle=\langle\bm{r},\bm{d}_{j}\rangle/\|\bm{d}_{j}\|_{2}$, together with $\bm{x}=\bm{D}\bm{z}=\bm{D}\,\bm{W}^{-1}\bm{W}\bm{z}=\widehat{\bm{D}}\,\widehat{\bm{z}}$. Since diagonal rescaling of the columns in $\bm{D}_{S}$ leaves the column span unchanged, the orthogonal projector onto $\mathrm{span}(\bm{D}_{S})$ is invariant to such rescaling.
Therefore, once the same index is selected, the least-squares updates utilize the same projector, and the residuals match at every step. ∎
Definition C.2 (ERC on normalized dictionary).
 [^7] We use $\mathrm{ERC}$ for the Tropp quantity whose threshold at $1$ is the classical exact recovery condition of Tropp [67].
@@
$\mathrm{ERC}(\widehat{\bm{D}};S)$ $:=\max_{j\in S^{c}}\bigl\|\,\widehat{\bm{D}}_{S}^{\dagger}\,\widehat{\bm{d}}_{j}\,\bigr\|_{1},$ (23)
$\mathrm{ERC}(\widehat{\bm{D}};S\,|\,T)$ $:=\max_{j\in T\setminus S}\bigl\|\,\widehat{\bm{D}}_{S}^{\dagger}\,\widehat{\bm{d}}_{j}\,\bigr\|_{1},$ (24)
for any $T\supseteq S$, where $\|\cdot\|_{1}$ denotes the vector $\ell_{1}$ norm. Here $S^{c}:=[k]\setminus S$, so $T\setminus S=T\cap S^{c}\subseteq S^{c}$; thus $\mathrm{ERC}(\widehat{\bm{D}};S\,|\,T)$ is the same maximum as $\mathrm{ERC}(\widehat{\bm{D}};S)$, but over the smaller index set $T\setminus S$.
Lemma C.3 (Monotone ERC improvement under subtree restriction).
@@
By definition, $\mathrm{ERC}(\widehat{\bm{D}};S\,|\,T)=\max_{j\in T\setminus S}\|\widehat{\bm{D}}_{S}^{\dagger}\widehat{\bm{d}}_{j}\|_{1}$. Since $\widehat{\bm{D}}_{S}^{\dagger}$ is fixed, shrinking $T$ only restricts the index set over which this maximum is taken, so the value cannot increase. ∎
Lemma C.4 (ERC threshold implies restricted OMP success).
Under the assumptions of Lemma C.3, if $\mathrm{ERC}(\widehat{\bm{D}};S\,|\,T_{L})<1$, then OMP run on $\bm{D}$ with the normalized selection rule
@@
Lemma C.1 shows that the normalized-selection rule on $\bm{D}$ matches standard OMP on $\widehat{\bm{D}}$. The classical noiseless ERC theorem of Tropp [67] applied to the restricted dictionary $\widehat{\bm{D}}_{T_{L}}$ then yields exact support recovery in $s$ iterations whenever $\max_{j\in T_{L}\setminus S}\|\widehat{\bm{D}}_{S}^{\dagger}\widehat{\bm{d}}_{j}\|_{1}<1$, equivalently whenever $\mathrm{ERC}(\widehat{\bm{D}};S\,|\,T_{L})<1$. ∎
C.5 Proof of Proposition 4.1
@@
If the maximizer(s) in $\mathrm{ERC}(\widehat{\bm{D}};S)$ is not in $T_{L}$, pruning them ensures $\mathrm{ERC}(\widehat{\bm{D}};S\,|\,T_{L})<\mathrm{ERC}(\widehat{\bm{D}};S)$, so the restricted value can fall below $1$ while the global one remains at least $1$. Whenever this happens, Lemma C.4 certifies exact recovery for Hierarchical OMP on $T_{L}$, whereas the ERC test for OMP on the full dictionary fails.
Appendix D Step-by-step construction of a Hierarchical Concept Embedding